6x-3y=-1 (1,-6) find the parallel line
Answer by mananth(16949)

6 x + -3 y = -1
Find the slope of this line
make y the subject
-3 y = -6 x + -1
Divide by -3
y = 2 x + 0.33
Compare this equation with y=mx+b
slope m = 2 2
The slope of a line parallel to the above line will be the same
The slope of the required line will be 2
m= 2 ,point ( 1 , -6 )
Find b by plugging the values of m & the point in y=mx+b
-6 = 2.00 + b
b= -8
m= 2
Plug value of b in y = mx +b
The required equation is y = 2 x -8
y = 2 x -8
m.ananth@hotmail.ca